OpenStudy (anonymous):

Convert 20 g/m.s to lbf.hr/ft^2 using the unit factor method.

OpenStudy (eyust707):

g/m.s ??? lbf.hr/ft^2?? not sure what those units are

OpenStudy (anonymous):

grams/meter.second; pound force.hour/feet^2

OpenStudy (anonymous):

\[20 \frac{g}{m.s}*\frac{1 lb_m}{454g}*\frac{0.3048m}{1ft}*\frac{1 lb_f}{32.174 \frac{ft.lb_m}{s^2}}*\frac{3600s}{1hr}=1.5024\frac{lb_f .hr}{ft^2}\]

OpenStudy (anonymous):

Are you sure that's right mukushla?

OpenStudy (anonymous):

i made a little mistake... \[20 \frac{g}{m.s}*\frac{1 lb_m}{454g}*\frac{0.3048m}{1ft}*\frac{1 lb_f}{32.174 \frac{ft.lb_m}{s^2}}*\frac{1 hr}{3600s}=1.159*10^{-7}\frac{lb_f .hr}{ft^2}\] plz dont tell me this is wrong again..........lol,,:D